OpenStudy (blackops2_143):

I will give a medal: Solve. -5/4n=40

OpenStudy (briannaontimeforreal):

i dont get it

pooja195 (pooja195):

Mutliply both sides by the reciprocal -5/4 ---> 4/-5 \[\huge~\rm~\bf~ \frac{ -5 }{ 4 } \times \frac{ 4 }{ -5 }n=40 \times \frac{ -5 }{ 4 }\]

OpenStudy (briannaontimeforreal):

oh i get it now

OpenStudy (blackops2_143):

Ok thx